OpenStudy (anonymous):

For questions 1–2, what is the solution of the system? Use a graph. 1. y = –x + 2 y = 3x – 1 (1 point) 2. –4x + 3y = –12 –2x + 3y = –18 (1 point)

OpenStudy (anonymous):

@zpupster would it help if i gave u the options?

OpenStudy (zpupster):

where the lines intersect is the solution. we can take 1) and figure it out algebraically. y = 3x – 1 y = –x + 2 subtract this equation that changes the sign in each term y = 3x – 1 -y = +x -2 0 = 4x - 3 ad 3 to both sides 3 = 4x and divide both sides by 4 3/4 = x so this is the answer for x now plug in x or 3/4 into one of the equations y = 3(3/4) – 1 y = 9/4 - 4/4---4/4 = 1 here y= 5/4 or 1 1/4 or 1.25 so are solution is (.75.1.25) you can see this on the graph now do no 2 https://www.desmos.com/calculator

OpenStudy (anonymous):

thank you sooo much @zpupster

OpenStudy (zpupster):

yw

OpenStudy (anonymous):

@zpupster how do i give medals?

OpenStudy (zpupster):

you already did thanks just choose best response

OpenStudy (anonymous):

ok cool @zpupster